OpenStudy (anonymous):

midas has 189 silver coins and 441 gold coins, he wants to place the gold coins and silver coins in stacks so that there are the same # of coins in each stack. What is the greatest # of coins that he can place in each stack?

OpenStudy (anonymous):

huh?

OpenStudy (dumbcow):

look for greatest factor

OpenStudy (anonymous):

a

OpenStudy (dumbcow):

189 = 3*63

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

189 = 3 * 3 *3 * 7 441 = 3 * 3 * 7 * 7 Greatest amount is 63